The perimeter of a rectangle is 68 cm. If the width is increased by 2cm. and twice th length is decreased by 10cm, the new perimeter is 100 cm. What are the dimensions of the original rectangle?
Answer by mananth(16949)

let width be x
length = y
2(x+y)=68
x+y=34.................1
..
x+2 new width
new length = 2y-10
2(x+2)+2(2y-10)=100
2x+4+4y-20=100
2x+4y-16=100
2x+4y=100+16
2x+4y=116
x+2y=58................2
subtract the equations
x+y-x-2y=34-58
-y=-24
y=24cm the width.
plug the value of y in equation 1
x+24=34
x=10cm the length